Roberto investment some money at 6% and then invested $5,000 more than twice this amount at 10% his total annual income from two investments was $3100. how much was invested at 10%
Answer by lwsshak3(11628)

Roberto investment some money at 6% and then invested $5,000 more than twice this amount at 10% his total annual income from two investments was $3100. how much was invested at 10%
***
let x=amt invested at 6%
2x+5000=amt invested at 10%
.06x+.10(2x+5000)=3100
.06x+.20x+500=3100
.26x=2600
x=10000
2x+5000=20000+5000=25000
how much was invested at 10%: $25,000